Jet Spectroscopy of the D_{1} 2^{2}B_{1} – D_{0} 1^{2}B_{1} Electronic Transition of the p-Cyanobenzyl Radical

\underline{Masaru Fukushima}, Advanced Technology Research Center, Mitsubishi Heavy Industries, LTD., Sachiura, Kanazawa, Yokohama 236, Japan; Kinichi Obi, Department of Chemistry, Tokyo Institute of Technology, Ohokayama, Meguro, Tokyo 152, Japan.

We have generated the p-cyanobenzyl radical in supersonic free expansion, and measured the laser induced fluorescence ( LIF ) spectra in green-blue region; the vibrationally and rotationally resolved LIF excitation spectra and the LIF dispersed spectra from the single vibronic levels ( SVL ). On the basis of the vibronic structures in the SVL dispersed spectra, the lowest energy band at 20735 cm^{−1} with the strongest intensity in the excitation spectrum has been assigned to the 0_{0}^{0} band of the visible spectrum. Based on the band type of the 0_{0}^{0} band, a-type, determined from the rotationally resolved LIF excitation spectrum, we have definitely assigned the visible band to the D_{1} 2^{2}B_{1} – D_{0} 1^{2}B_{1} electronic transition. We have found on the grounds of the vibrational analysis of the dispersed spectra, that the vibronic structure of the 2^{2}B_{1} – 1^{2}B_{1} electronic transition of the benzyl type was characterized by totally symmetric fundamental modes, 1, 8a, and 9a, with the frequencies higher than 800 cm^{−1}, while those of the 1^{2}A_{2} – 1^{2}B_{1} electronic transition are non-totally symmetric modes, 6b and 8b, and totally symmetric modes, 6a and 1, with the frequencies lower than 800 cm^{−1} ^{a}.

aM. Fukushima, K. Obi, Chem. Phys. Lett. 248, 269 (1996)